One hundred dollars is invested at 12% interest compounded annually. Determine how much the investment is worth after 1 year.

Answer by mananth(16949)

Principal P = 100
Amount= A
years=n 1
compounded 1 times a year t
Simple interest & compound Interest for first year are same
Rate = 12.00 0.12
Amount =P*((n+r)/n)^n*t

Amount =100 *( 1 + 0.12 )^ 1 * 1
Amount =100*(1+0.12)^1
100*(1.12)^ 1
Amount = 112